# Optimal Tests of the Composite Null Hypothesis Arising in Mediation Analysis

### v2

@@ 7.1 Single mediation hypothesis testing
To compare the finite sample performances of our proposed tests with those of the delta method, joint significance, and van Garderen and van Giersbergen, 2022a (henceforth vGvG) tests, we performed a simulation study in which we sampled from independent $t$ distributions with noncentrality parameters $(\delta_{x},\delta_{y})$. We considered four scenarios: (a) We varied $\delta_{x}$ and $\delta_{y}$ jointly from 0 to 0.4. (b) We fixed $\delta_{y}$ to 0.2 and varied $\delta_{x}$ from 0 to 0.4. (c) Using a normal approximation, we fixed $\delta_{y}$ to 0 and varied $\delta_{x}$ from 0 to 0.4, thereby exploring empirical type 1 error across a range of parameter values within the composite null hypothesis space. (d) We used the same setting as in (c), but using a $t$-distribution with $n-1$ degrees of freedom approximation instead of the normal approximation. We drew 100,000 Monte Carlo samples with $n=50$ for each value of $\delta$, and applied each test of $H_{0}:``\delta_{x}\delta_{y}=0"$ against $H_{1}:``\delta_{x}\delta_{y}\neq 0"$ to each sample. In particular, we used three versions of the Bayes risk optimal test: with the 0-1 loss function, with the quadratic loss function, and the constrained test with quadratic loss function. Likewise, we used three version of the minimax optimal test: the standard version, the test based on the corresponding $p$-value defined in the Supplementary Materials, and a truncated minimax optimal test. In both the constrained and truncated tests, we used $d=0.1$ in order to make a fair comparison with the vGvG test, whose rejection region is 0.1 away from the null hypothesis space. The Monte Carlo power estimates are displayed in the plots in Figure 4.
@@
Figure 4: Monte Carlo rejection probabilities of the following: Bayes risk optimal with 0-1 loss (BRO 0-1), Bayes risk optimal with quadratic loss (BRO quad), constrained Bayes risk optimal with quadratic loss (BRO q-cnst), delta method (Delta), joint significance (JS), minimax optimal (MMO), $p$-value-based test of the minimax optimal test (MMO pval), truncated minimax optimal (MMO trunc), and vGvG. (a) $\delta_{x}=\delta_{y}$ varying from 0 to 0.4. (b) $\delta_{y}$ fixed at 0.2, $\delta_{x}$ from 0 to 0.4. (c) $\delta_{y}$ fixed at 0, $\delta_{x}$ from 0 to 0.4 with normality-based tests. (d) $\delta_{y}$ fixed at 0, $\delta_{x}$ from 0 to 0.4. with $t$-distribution-based tests.
@@
In scenario (a), apart from the minimax optimal $p$-value-based test, all of the minimax optimal and Bayes risk optimal tests and the vGvG test have very close to nominal type 1 error at $(\delta_{x},\delta_{y})=(0,0)$. All of these tests greatly outperform the delta method and joint significance tests in terms of power for smaller values of $\delta_{x}=\delta_{y}$, with the minimax optimal $p$-value-based test having power somewhere in between, but converging more quickly to the other tests. The delta method test remains very conservative over the entire range of parameter values. All of the tests except for the delta method test begin to converge in power closer to 0.4. The minimax optimal (apart from the $p$-value based test), Bayes risk optimal, and vGvG tests perform very similarly over the range of parameter values, with the quadratic loss versions of the Bayes risk optimal test trading off some power loss in the 0.1–0.2 range for a slight improvement in power in the 0.3–0.4 range, demonstrating the greater emphasis the quadratic loss places on larger alternatives. The truncated minimax optimal test begins slightly conservative relative to 0.05, but catches up with the most powerful tests fairly quickly. Despite not being similar tests, the Bayes risk optimal and vGvG tests suffer very little in terms of power near the least favorable distribution at $(\delta_{x},\delta_{y})=(0,0)$.
@@
In scenario (b), the trends are largely the same as in scenario (a). The main difference is that the quadratic loss versions of the Bayes risk optimal test are slightly conservative with respect to 0.05 under the null, albeit less conservative than the minimax optimal $p$-value-based test. Once again, these tests trade off some power loss for smaller values of $\delta_{x}$ (values less than about 0.25) for slightly improved power for larger values of $\delta_{x}$ (from about 0.35–0.55). Since $\delta_{y}$ is fixed at 0.2, power for all tests appears to plateau a little under 0.3. In scenarios (c) and (d), we see that all tests approximately preserve type 1 error; however, there is some anti-conservative behavior in scenario (c) due to the normal approximation being a bit too concentrated for a sample size of 50. This is corrected with the use of the $t$-distribution approximation, which yields type 1 error much closer to 0.05 across the entire range of the null hypothesis parameter space we consider. In both cases, we see the performances of all tests except the delta method to converge by around $\delta_{x}=0.55$.
@@ 7.2 Large-scale mediation hypothesis testing
Table 1: Mean (standard deviation) of the ratio of the actual FPR to the nominal FPR using the joint significance, JT-comp (Huang,, 2019), HDMT (Dai et al.,, 2020), Sobel-comp (Du et al.,, 2023), DACT (Liu et al.,, 2022), and minimax optimal tests.
@@ 8.1 DCTRS Data Analysis
We applied our methodology to data from the Database of Cognitive Training and Remediation Studies (DCTRS), which consists of data from several randomized trials testing the efficacy of cognitive remediation therapy in patients with schizophrenia. Cognitive remediation therapy targets patients’ cognitive outcomes with the long-term goal of cognitive gains translating into improvements in more distal outcomes, such as functioning and quality of life. We used data from three trials described in Wykes et al., (1999), Wykes et al., 2007a , and Wykes et al., 2007b , consisting of 128 patients with complete treatment, mediator, and outcome data. In each study, patients were randomized to a cognitive remediation therapy arm or a control arm. There are a number of issues (e.g., potential unobserved mediator-outcome confounding, variable follow-up time, and the fact that the data come from studies with different protocols) that we do not entirely account for in this analysis, and so our results should be viewed as purely illustrative of the proposed methodology.
@@
Our exposure of interest, $A$, was an indicator of whether the patient was assigned to the treatment arm; our outcome of interest, $Y$, was the patient’s Rosenberg self-esteem scale score at follow-up, and the potential mediator of interest, $M$, was the patient’s Wechsler Adult Intelligence Scale (WAIS) working memory digit span test at their end of treatment, which measures both working memory and attention. The mean self-esteem scale score was 33.6 with sample range from 17 to 50. Since the studies were all randomized trials, we only need to assume that variables included in $\bm{C}$ and $A$ are sufficient to control for confounding of the effect of $M$ on $Y$. We let $\bm{C}$ consist of the patients’ sex, race, education, which study they participated in, and their self-esteem scale and working memory digit span test scores measured at baseline. We provide a more detailed discussion about the variable choices and modeling assumptions in the Supplementary Materials. Previously, Wykes et al., 2007b found cognitive remediation to improve working memory digit span test scores; Wykes et al., (1999) found cognitive remediation to improve self-esteem as measured by the Rosenberg self-esteem scale. We are testing whether there is an effect of cognitive remediation on patients’ self-esteem that is mediated by its effect on working memory and attention.
@@
We fit models (2) and (4) using ordinary least squares, and estimated the natural indirect effect by the corresponding estimator given in Section 2. We then applied the three versions of the minimax optimal test and the three versions of the Bayes risk optimal test from Section 7, and compared these with the joint significance, delta method, and vGvG tests, all at level $\alpha=0.05$, using the t-distribution approximation. The estimated natural indirect effect is 0.29, which is interpreted as cognitive remediation therapy modestly improving the self-esteem scale score by an average of 0.29 (0.04 standard deviations of the self-esteem scale score) through its effect on working memory and attention as measured by the digit span test.
@@
The test statistic is plotted on the rejection regions of the minimax optimal and Bayes risk optimal tests in Figures S4–S8 of the Supplementary Materials.
Here we observe the improved power of the different versions of the minimax optimal test to reject the null hypothesis that the NIE is zero in favor of the alternative that it is nonzero. The interpretation for any of these tests is that if the true NIE were in fact zero, then the event we observed would have been unlikely to have occurred in the sense that the test would have rejected with at most 5% probability. Since these tests disagree with the joint significance test, they lack the stronger interpretation that the probability of observing an event as or more extreme would be at most 5% if the true NIE were zero. Having established that what we have observed is “unlikely” in the former sense, one may proceed by examining whether the estimated effect size is meaningful. An effect size of 0.04 standard deviations of the self-esteem scale score may not be deemed relevant for understanding the mechanism by which cognitive remediation therapy affects self-esteem. However, had the estimate been larger, one might conclude that there was a meaningful indirect effect through working memory, and that such an estimate would have been unlikely to have been observed if there were no true underlying indirect effect. In this case, neither of the traditional tests nor any of the Bayes risk optimal tests reject, but the vGvG test does reject, agreeing with the minimax optimal tests. While we have conducted nine tests for purposes of illustration, in practice one should decide a priori which test to use based on how they wish to navigate the power–interpretability trade-off to avoid data dredging.
@@ 8.2 Normative Aging Study Analysis
Liu et al., (2022) performed large-scale hypothesis testing on the Normative Aging Study (NAS) to test the mediated effect of smoking status on forced expiratory flow at 25%–75% of the Forced Expiratory Vital capacity, a measure of lung function, through 484,613 DNA methylation CpG cites among men in Eastern Massachusetts. We apply our proposed methodology to this same data and compare results. See Liu et al., (2022) for more details regarding this data. Based on the minimax optimal test with a Bonferroni correction controlling FWER at 0.05, we detect significant mediated effects through sites cg03636183, cg05575921, cg06126421, and cg21566642. The conservative version of the Benjamini–Hochberg correction controlling FDR at 0.05 additionally detected a mediated effect through the site cg05951221. The Bonferroni procedure ran locally on a MacBook Pro with 2.3 GHz Intel Core i5 in 7.3 seconds; the Benjamini–Hochberg procedure in 18.9 seconds. Liu et al., (2022) notes that each of these sites has previously been found to be related to smoking status and/or lung cancer risk. In comparison, the FDR-adjusted joint significance test identified all of these except cg05951221, while the FDR-adjusted DACT method of Liu et al., (2022) found all of these to be significant as well as 14 others, the latter of which they did not make any connections to previous findings in the literature.
@@ 9 DISCUSSION
We have proposed two novel classes of tests for the composite null hypothesis $H_{0}:``\delta_{x}\delta_{y}=0"$ against $H_{1}:``\delta_{x}\delta_{y}\neq 0"$, which arises commonly in the context of mediation analysis. This is a challenging inferential task due to the non-uniform asymptotics of univariate test statistics of the product of coefficients, as well as the nonstandard structure of the null hypothesis space. We have constructed these tests to be both optimal in a decision theoretic sense, and to preserve type 1 error uniformly over the null hypothesis space—exactly so on both counts in the case of the minimax optimal test, and approximately so in the case of the Bayes risk optimal test. We have described procedures for carrying out large-scale hypothesis testing of many product-of-coefficient hypotheses, controlling for both familywise error rate and false discovery rate. We have also considered some shortcomings of these tests in terms of interpretability, and have developed modifications to remedy some of these shortcomings. Each of these comes at some cost of power, and we have illustrated a fundamental trade-off between the objectives of power and interpretability. Lastly, we have provided an R package to implement the minimax optimal test.
@@
It is natural to ask which of the proposed tests should be used in practice. This is highly subjective, and depends on the criteria by which one judges hypothesis tests, as well as how much value one places on power in different parts of the alternative hypothesis space. We have introduced tests that improve on power of traditional tests, but not all of which have rejection regions bounded away from the null hypothesis space or that are monotonic in $\alpha$. If one is only concerned with maximizing power subject to the constraint of uniformly preserving type 1 error, then these tests will be desirable. If one is concerned about the possibility of rejecting with small values of the test statistic, then one may employ a truncated or constrained version of these tests. If one finds the nonmonotocity property inadmissible, then one can use the test based on the $p$-value corresponding to the minimax optimal test. If one is concerned about both, then one can use a truncated version of the $p$-value based test. The minimax optimal test has the advantage of being a closed form exact solution to its corresponding optimization problem as well as preserving type 1 error exactly in the limit. The minimax optimal test also can be generated almost instantaneously for all unit fraction values of $\alpha$, and the generalized test can be generated for other values of $\alpha$, whereas for the Bayes risk optimal test, the optimization step to generate the rejection region must be performed each time a test for a new value of $\alpha$ is desired. $p$-values are readily obtained for the minimax optimal test. Lastly, the minimax optimal test is a strictly deterministic test, whereas the Bayes risk optimal test contains a few cells in the rejection region that reject with non-degenerate probability. If a non-random test is desired, the latter can be converted to a slightly more conservative non-random test.
There are a number of important future research directions stemming from this work. There are many inferential problems that face similar issues with non-uniform convergence. Examples commonly arise in partial identification bounds, where there are often minima and maxima of estimates involved, which exhibit similar asymptotic behavior. The approaches proposed in this article can likely be adapted to such settings. Under different models, mediated effects can take the form of sums of products of coefficients. This considerably more complex setting will be important to address to allow for less restrictive models when performing inference on mediated effects. Lastly, we plan to further develop the tests for products of more than two coefficients discussed in the Supplementary Materials.

### v3

@@ 7.1 Single mediation hypothesis testing
To compare the finite sample performances of our proposed tests with those of the others we have discussed, we performed a simulation study in which we sampled from independent $t$ distributions with noncentrality parameters $\bm{\delta}$. We considered four scenarios: (a) We varied $\delta_{x}$ and $\delta_{y}$ jointly from 0 to 0.4. (b) We fixed $\delta_{y}$ to 0.2 and varied $\delta_{x}$ from 0 to 0.4. (c) Using a normal approximation, we fixed $\delta_{y}$ to 0 and varied $\delta_{x}$ from 0 to 0.4, thereby exploring empirical type 1 error across a range of parameter values within the composite null hypothesis space. (d) We used the same setting as in (c), but using a $t$-distribution with $(n-1)$ degrees of freedom approximation instead of the normal approximation. We drew 100,000 Monte Carlo samples with $n=50$ for each value of $\delta_{x}$, and applied each test of $H_{0}:``\delta_{x}\delta_{y}=0"$ against $H_{1}:``\delta_{x}\delta_{y}\neq 0"$ to each sample. In particular, we used three versions of the Bayes risk optimal test: the tests with either the 0-1 loss or quadratic loss function, and the constrained test with quadratic loss function. Likewise, we used three version of the minimax optimal test: the standard version, the test based on the corresponding $p$-value defined in the Supplementary Materials, and a truncated minimax optimal test. In both the constrained and truncated tests, we used $d=0.1$ in order to make a fair comparison with the van Garderen and van Giersbergen, (2024) (henceforth vGvG) test. The Monte Carlo power estimates are displayed in the plots in Figure 4.
@@
Figure 4: Monte Carlo rejection probabilities of the following: Bayes risk optimal with 0-1 loss (BRO 0-1), Bayes risk optimal with quadratic loss (BRO quad), constrained Bayes risk optimal with quadratic loss (BRO q-cnst), delta method (Delta), joint significance (JS), minimax optimal (MMO), $p$-value-based test of the minimax optimal test (MMO pval), truncated minimax optimal (MMO trunc), and vGvG. (a) $\delta_{x}=\delta_{y}$ varying from 0 to 0.4. (b) $\delta_{y}$ fixed at 0.2, $\delta_{x}$ from 0 to 0.6. (c) $\delta_{y}$ fixed at 0, $\delta_{x}$ from 0 to 0.8 with normality-based tests. (d) $\delta_{y}$ fixed at 0, $\delta_{x}$ from 0 to 0.8 with $t$-distribution-based tests. In panel d, the BRO quad and BRO q-cnst lines are identical.
@@
In scenario (a), apart from the minimax optimal $p$-value-based test, all of the minimax optimal and Bayes risk optimal tests and the vGvG test have very close to nominal type 1 error at $\bm{\delta}=(0,0)^{\top}$. All of these tests greatly outperform the delta method and joint significance tests in terms of power for smaller values of $\delta_{x}=\delta_{y}$, with the minimax optimal $p$-value-based test having power somewhere in between, but converging more quickly to the other tests. The delta method test remains very conservative over the entire range of parameter values. The minimax optimal (apart from the $p$-value based test), Bayes risk optimal, and vGvG tests perform very similarly over the range of parameter values, with the quadratic loss versions of the Bayes risk optimal test trading off some power loss in the 0.1–0.2 range for a slight improvement in power in the 0.3–0.4 range, demonstrating the greater emphasis the quadratic loss places on larger alternatives. The truncated minimax optimal test begins slightly conservative relative to 0.05, but catches up with the most powerful tests quickly. Despite not being similar tests, the Bayes risk optimal and vGvG tests suffer very little in terms of power near the least favorable distribution at $\bm{\delta}=(0,0)^{\top}$.
@@
In scenario (b), the trends are largely the same as in scenario (a). The main difference is that the quadratic loss versions of the Bayes risk optimal test are slightly conservative under the null, albeit less conservative than the minimax optimal $p$-value-based test. These tests trade off some power loss for smaller values of $\delta_{x}$ ($\lesssim$0.25) for slightly improved power for larger values of $\delta_{x}$ ($\approx$0.35–0.55). Since $\delta_{y}$ is fixed at 0.2, power for all tests appears to plateau a little under 0.3. In scenarios (c) and (d), we see that all tests approximately preserve type 1 error; however, there is some anti-conservative behavior in scenario (c) due to the normal approximation being a bit too concentrated for a sample size of 50. This is corrected with the use of the $t$-distribution approximation, which yields type 1 error much closer to 0.05 across the null hypothesis parameter space. In both cases, we see the performances of all tests except the delta method to converge by around $\delta_{x}=0.55$.
@@ 7.2 Large-scale mediation hypothesis testing
Table 1: Mean (standard deviation) of the ratio of the actual FPR to the nominal FPR using the joint significance, JT-comp (Huang,, 2019), HDMT (Dai et al.,, 2022), Sobel-comp (Du et al.,, 2023), DACT (Liu et al.,, 2022), and minimax optimal tests.
@@ 8.1 DCTRS Data Analysis
We applied our methodology to data from the Database of Cognitive Training and Remediation Studies (DCTRS), which consists of data from several randomized trials testing the efficacy of cognitive remediation therapy in patients with schizophrenia. Cognitive remediation therapy targets patients’ cognitive outcomes with the long-term goal of cognitive gains translating into improvements in more distal outcomes like functioning and quality of life. We used data from three trials described in Wykes et al., (1999), Wykes et al., 2007a , and Wykes et al., 2007b consisting of 128 patients with complete treatment, mediator, and outcome data. In each study, patients were randomized to a cognitive remediation therapy arm or a control arm. There are a number of issues (e.g., potential unobserved mediator-outcome confounding, variable follow-up time, and the fact that the data come from studies with different protocols) that we do not entirely account for in this analysis, and so our results should be viewed as purely illustrative of the proposed methodology.
@@
Our exposure of interest, $A$, was an indicator of whether the patient was assigned to the treatment arm; our outcome of interest, $Y$, was the patient’s Rosenberg self-esteem scale score at follow-up, and the potential mediator of interest, $M$, was the patient’s Wechsler Adult Intelligence Scale (WAIS) working memory digit span test at their end of treatment, which measures both working memory and attention. The mean self-esteem scale score was 33.6 with sample range from 17 to 50. Since the studies were all randomized trials, we only need to assume that variables included in $\bm{C}$ and $A$ are sufficient to control for confounding of the effect of $M$ on $Y$. We let $\bm{C}$ consist of the patients’ sex, race, education, which study they participated in, and their self-esteem scale and working memory digit span test scores measured at baseline. We provide more detailed discussion about the variable choices and modeling assumptions as well as unadjusted and more comprehensively adjusted analyses in the Supplementary Materials. All tests agree in these additional analyses. The relationship between covariate adjustment and agreement of the various tests is complex. Adjusting for more covariates can influence the underlying test statistics in either direction, and doing so does not consistently make the various tests more or less likely to agree. Previously, Wykes et al., 2007b found cognitive remediation to improve working memory digit span test scores; Wykes et al., (1999) found cognitive remediation to improve Rosenberg self-esteem scale scores. We are testing whether there is an effect of cognitive remediation on patients’ self-esteem that is mediated by its effect on working memory and attention.
@@
We fit models (2) and (4) using ordinary least squares, and estimated the natural indirect effect by the corresponding estimator given in Section 2. We then applied the three versions of the minimax optimal test and the three versions of the Bayes risk optimal test from Section 7, and compared these with the joint significance, delta method, and vGvG tests, all at level $\alpha=0.05$, using the t-distribution approximation. The estimated total effect is 0.70, which is interpreted as cognitive remediation therapy modestly improving the self-esteem scale score by an average of 0.70 (0.10 standard deviations). The estimated natural indirect effect is 0.29, which is interpreted as cognitive remediation therapy modestly improving self-esteem by an average of 0.29 (0.04 standard deviations) through its effect on working memory and attention as measured by the digit span test.
@@
The estimated proportion mediated is 41.0%. The test statistic is plotted on the rejection regions of the minimax optimal and Bayes risk optimal tests in Figures S4–S8 of the Supplementary Materials.
Hypothesis test results are presented in Table 2, where we observe the improved power of the different versions of the minimax optimal test to reject the null hypothesis that the NIE is zero in favor of the alternative that it is nonzero. The interpretation for any of these tests is that if the true NIE were in fact zero, then the event we observed would have been unlikely to have occurred in the sense that the test would have rejected with at most 5% probability. Since these tests disagree with the joint significance test, they lack the stronger interpretation that the probability of observing an event as or more extreme would be at most 5% if the true NIE were zero. Having established that what we observed is “unlikely” in the former sense, one may proceed by examining whether the estimated effect size is meaningful. An effect size of 0.04 standard deviations of the self-esteem scale score may not be deemed relevant for understanding the mechanism by which cognitive remediation therapy affects self-esteem. However, had the estimate been larger, one might conclude that there was a meaningful indirect effect through working memory, and that such an estimate would have been unlikely to have been observed if there were no true underlying indirect effect. Alternatively, one might consider a proportion mediated exceeding 10%, say, to be meaningful, and would then conclude that the indirect effect through working memory indeed explains a meaningful amount of the total effect of cognitive remediation. However, as the total effect is not statistically significant in this case, the proportion mediated may be a less compelling criterion for evaluating practical significance. In this case, neither of the traditional tests nor any of the Bayes risk optimal tests reject, but the vGvG test does reject, agreeing with the minimax optimal tests. While we have conducted nine tests for purposes of illustration, in practice one should decide a priori which test to use based on how they wish to navigate the power–interpretability trade-off to avoid data dredging.
@@ 8.2 Normative Aging Study Analysis
Liu et al., (2022) performed large-scale hypothesis testing on the Normative Aging Study (NAS) to test the mediated effect of smoking status on forced expiratory flow at 25%–75% of the Forced Expiratory Vital capacity, a measure of lung function, through 484,613 DNA methylation CpG cites among men in Eastern Massachusetts. We apply our proposed methodology to this data and compare results. See Liu et al., (2022) for details regarding this data. Based on the minimax optimal test with a Bonferroni correction controlling FWER at 0.05, we detect significant mediated effects through sites cg03636183, cg05575921, cg06126421, and cg21566642. The conservative version of the Benjamini–Hochberg correction controlling FDR at 0.05 additionally detects a mediated effect through the site cg05951221. The Bonferroni procedure ran locally on a MacBook Pro with 2.3 GHz Intel Core i5 in 7.3 seconds; the Benjamini–Hochberg procedure in 18.9 seconds. Liu et al., (2022) note that each of these sites has previously been found to be related to smoking status and/or lung cancer risk. In comparison, the FDR-adjusted joint significance test identified all of these except cg05951221, while the FDR-adjusted DACT method of Liu et al., (2022) found all of these to be significant as well as 14 others, the latter of which do not appear to have been found to be linked with either smoking or lung cancer. Furthermore, in our simulation study the DACT was found to have an inflated proportion of false positives—over 22 times that of the nominal proportion when using the appropriate cutoff—whereas our minimax optimal test never identified false positives at a rate any higher than 1.35 times that of the nominal proportion. These considerations cast doubt on whether the additional 14 CpG sites identified by the DACT method but not by ours are true positives. Thus, the FDR-adjusted minimax optimal test identified more CpG sites than the standard joint significance or delta method tests while also preserving approximate FDR control as shown theoretically and in the simulation study.
@@ 9 DISCUSSION
We have proposed two novel classes of tests for the composite null hypothesis $H_{0}:``\delta_{x}\delta_{y}=0"$ against $H_{1}:``\delta_{x}\delta_{y}\neq 0"$, which arises commonly in the context of mediation analysis. This is a challenging inferential task due to the non-uniform asymptotics of univariate test statistics of the product of coefficients, as well as the nonstandard structure of the null hypothesis space. We have constructed these tests to be both optimal in a decision theoretic sense, and to preserve type 1 error uniformly over the null hypothesis space—exactly so on both counts in the case of the minimax optimal test, and approximately so in the case of the Bayes risk optimal test. We have described procedures for carrying out large-scale hypothesis testing of many product-of-coefficient hypotheses, controlling for both familywise error rate and false discovery rate. We have also considered some shortcomings of these tests in terms of interpretability, and have developed modifications to remedy some of these. Each of these comes at some cost of power, and we have illustrated a fundamental trade-off between the objectives of power and interpretability. Lastly, we have provided an R package to implement the proposed tests.
It is natural to ask which of the proposed tests should be used in practice. This is highly subjective, and depends on the criteria by which one judges hypothesis tests, as well as how much value one places on power in different parts of the alternative hypothesis space. We summarize the trade-offs with respect to power, scalability for large-scale hypothesis testing, and interpretability of the various tests discussed in this article in Table 3.
Table 3: Summary of properties of various tests
Hypothesis test Power Scalability to Monotonicity Information Bounded away
mult. testing in $\alpha$ coherence from null
Bayes risk optimal:
– 0-1 loss highest no can be no yes
– quad. loss highest no can be no no
– quad. loss, constr. higher no can be no yes
Minimax optimal:
– standard highest yes no no no
– truncated higher yes no no yes
$p$-value moderate yes yes no no
Delta method lowest yes yes yes yes
Joint significance lower yes yes yes yes
vGvG highest no yes no yes
@@
We have introduced tests that improve on the power of traditional tests, but not all of which have rejection regions bounded away from the null hypothesis space or that are monotonic in $\alpha$. If one is only concerned with maximizing power subject to the constraint of uniformly preserving type 1 error, then these tests will be desirable. If one is concerned about the possibility of rejecting with small values of the test statistic, then one may employ a truncated or constrained version of these tests. If one finds the nonmonotocity property inadmissible, then one can use the test based on the $p$-value corresponding to the minimax optimal test. If one is concerned about both, then one can use a truncated version of the $p$-value based test. The minimax optimal test has the advantage of being a closed form exact solution to its corresponding optimization problem as well as preserving type 1 error exactly in the limit. The minimax optimal test also can be generated almost instantaneously for all unit fraction values of $\alpha$, and the generalized test can be generated for other values of $\alpha$, whereas for the Bayes risk optimal test, the optimization step to generate the rejection region must be performed each time a test for a new value of $\alpha$ is desired. It is easy to obtain $p$-values for the minimax optimal test. Lastly, the minimax optimal test is a strictly deterministic test, whereas the Bayes risk optimal test contains a few cells in the rejection region where one rejects with non-degenerate probability. If a non-random test is desired, the latter can be converted to a slightly more conservative non-random test.
We close with some practical recommendations for implementation of our proposed methodology. We reiterate that which test to use is subjective and comes down to how the investigator thinks about interpretability and weighs the properties summarized in Table 3. One will often wish to avoid reporting significant findings that are not “practically relevant”, in which case we would recommend pre-specifying one’s criterion of “practical relevance”. This might be in terms of the proportion mediated, say if the objective is to quantify how much of a significant total effect is transmitted through the candidate mediator, or it could be in terms of the indirect effect estimate itself if this is considered to be the more relevant measure to threshold. Alternatively, if one wishes to take into account uncertainty, then one might prefer to threshold based on the test statistics, which would correspond to using the truncated minimax optimal or constrained Bayes risk optimal test.